OpenStudy (anonymous):

Solve, to one decimal place, using the quadratic formula. 2x^2 + 6x - 5 = 0, Show steps please. thaank you

OpenStudy (kc_kennylau):

\(x=\dfrac{-6\pm\sqrt{6^2-4\times2\times(-5)}}{2\times2}=\)your turn